# Cloud2BIM: An open-source automatic pipeline for efficient conversion of large-scale point clouds into IFC format

### v1

@@ 3.4 Identification of zones
The detection of zones is typically approached through three primary methods (Chen et al., 2024): (i) slice-based modeling, (ii) mesh simplification modeling, and (iii) floorplan-based modeling. Once the exact positions and geometries of walls are identified, Cloud2BIM circumvents these computationally intensive methods. Instead, the identify_zones function (Figure LABEL:fig:zones_flochart) analyzes wall geometry by first processing the axes of walls to ensure connectivity and proper alignment. Wall segments are split at intersections of these axes and extended to connect with adjacent axes within a distance specified by the snapping_distance parameter. Subsequently, segments parallel to the wall axes, representing the surfaces of the walls, are generated at a distance equal to half the width of the individual walls. These parallel segments are then extended to intersect fully, forming the boundaries of zones. The function ultimately creates closed polygons defined by the wall axes; each polygon serves as a candidate zone. The complete process of zone identification is illustrated in Figure 12.
Figure 11: Flowchart diagram illustrating the identify_zones function.
@@ 3.5 Identification of openings
Several methods for detecting openings have been developed over the years. Techniques for identifying door openings have typically been categorized into histogram-based approaches (Mahmoud et al., 2024; Chen, 2022) and those that utilize scanner trajectory data (Nikoohemat et al., 2018). Each method, however, comes with specific limitations. Density-based analysis can be sensitive to obstructions like furniture, affecting the histogram’s accuracy, while trajectory-based methods depend heavily on particular scanning technology. Due to these constraints, our approach utilizes a density analysis from histograms, as illustrated in the flowchart in Figure 13, where the function identify_openings is outlined.
Figure 13: Flowchart diagram illustrating the identify_openings function.
@@
The analysis begins by positioning the wall at the coordinate origin and rotating it into the $x$-$z$ plane. The floor plan position of the opening, primarily along the $x$-axis, is first determined. The histogram is then analyzed based on a threshold defined as a percentage of the tenth maximum value recorded to identify the start and end of gaps or low-density areas. The tenth maximum was selected based on observations of the histograms and found appropriate to avoid extreme maxima that seldom occur in the histogram; in some walls there were about 1–3 extreme maxima so that the tenth maximum provides a safe reserve, but the number of selected maximum can be changed. When matching openings are detected, this procedure is repeated for the $z$-axis histogram. Figure 14 presents a side view of the wall, displaying histograms for the $x$ and $z$ axes with marked door openings in red. The effectiveness of this analysis is improved by the prior dilution of the point cloud to the minimum distance, as discussed in Section 2. A more uniform point distribution minimizes the risk of misjudging the thresholds for empty space in the histograms, thereby improving the accuracy of opening detection.
@@ 3.6 Creation of IFC output file
The industry foundation classes (IFC) file format is primarily designed to facilitate data transfer among various participants in the modeling process, allowing the sharing of the building information model and its associated data. Utilizing IFC as the final format to transfer data from point clouds to BIM is a well-established approach, as evidenced by similar implementations by Martens and Blankenbach (2023), which included the reconstruction of slabs, floors, and walls. There are numerous scan-to-BIM methodologies discussed in the literature, such as those by Adán et al. (2018) and Truong-Hong and Lindenbergh (2022); however, these approaches typically do not support direct export to IFC or any standardized BIM format. The capability to export directly to IFC, as implemented in our Cloud2BIM tool, facilitates seamless integration with a wide range of BIM software platforms. This compatibility significantly enhances the utility of the generated models for architectural planning, structural analysis, and lifecycle management, ensuring that stakeholders can leverage the full spectrum of BIM functionalities for efficient project execution and maintenance.
Our output of Cloud2BIM is generated in the IFC 4 ADD2 TC1 format using the IfcOpenShell library for Python (IfcOpenShell contributors, 2023). This process includes specifying all necessary attributes for each IFC class. We chose the DesignTransferView_V1.0 Model View Definition to ensure that our models are compatible with and can be further edited in BIM software platforms such as ArchiCAD [^3] https://graphisoft.com/solutions/archicad, Revit [^4] https://www.autodesk.com/products/revit, and Allplan [^5] https://www.allplan.com.
@@ 4.1 Comparative analysis and benchmarking of algorithms
Multi-story
Slabs
Oblique wall axes
Wall surfaces
Volumetric walls
Volumetric spaces
Openings
Stairs
Content elements
IFC output
Open-source
Wang et al. Wang et al. (2015)
Macher et al. Macher et al. (2017)
Jung et al. Jung et al. (2018)
Shi et al. Shi et al. (2019)
Ochmann et al. Ochmann et al. (2019)
Bassier and Vergauwen Bassier and Vergauwen (2020)
Romero-Jarén and Arranz Romero-Jarén and Arranz (2021)
Gen and Gentes Chen and Gentes (2021)
Martens and Blankenbach Martens and Blankenbach (2023)
Mahmoud et al. Mahmoud et al. (2024)
Mehranfar et al. Mehranfar et al. (2024a)
@@
In reviewing the contributions of others, our discussion emphasizes the distinction in methodologies, scope, and outcomes. We highlight how different authors have approached the segmentation of point clouds and the geometric reconstruction of architectural elements, pointing out the strengths and limitations of their approaches compared to ours. For instance, Bassier and Vergauwen (2020) focuses solely on walls, creating geometry in Revit through a Rhino plug-in and then exporting to IFC, including zones. Our method supports direct IFC export from Python, reducing dependence on commercial software. In addition, their approach, like some others, lacks the ability to detect and model doors and windows, often resulting in a lower level of detail. This study though, unlike the other, allows
@@
The work of Mahmoud et al. (2024) incorporates advanced deep learning techniques for semantic segmentation, followed by RANSAC for line extraction to identify architectural features. While it provides detailed segmentation of walls, floors, and ceilings, the geometry export is limited to an Excel format, which is then used by a Dynamo add-in for Revit, constraining the output to specific commercial software. Similarly, the work by Mehranfar et al. (2024b) relies on the Revit API for 3D geometry and employs the PointTransformer network with DBScan clustering. However, this approach is sensitive to auxiliary parameter settings and does not enhance functionality beyond our method. Furthermore, it only supports single-floor point clouds, limiting its utility for multi-storey buildings.
@@
Martens and Blankenbach (2023) presented VOX2BIM+, a robust approach by analyzing voxels to improve noise resistance and accuracy in the volumetric reconstruction of spaces and structural elements. Unlike our approach, which utilizes direct point cloud processing, their method incorporates several advanced techniques, including region growing and morphological skeletons for wall detection. However, their code remains proprietary, and their method does not support window or door geometry extraction.
@@
Similarly, Wang et al. (2015) and Ochmann et al. (2019) have made significant contributions to the field, focusing on energy simulation models to detailed wall segmentation. These studies typically rely on simplified models or specific algorithms like RANSAC, which may only capture some architectural details such as doors and windows, a capability our software aims to address comprehensively. While sufficient for their goals, the detected zones that are clustered according to their purpose and not published to IFC cannot be further edited and exported to other software.
@@ 4.2 Precision analysis
This section evaluates the precision of our segmentation, classification, and reconstruction algorithms across different datasets as detailed in Section 2.
@@ 4.2.1 ETH Zürich synth3
The ETH Zürich synth3 dataset was selected to showcase the performance of our volumetric method in wall segmentation within a 2.5-dimensional Atlanta-world system. This dataset is frequently used to benchmark the success of interior wall segmentation (Ochmann et al., 2019; Bassier and Vergauwen, 2020; Shi et al., 2019). Although it lacks an exterior scan, this limitation mirrors real-life scenarios where facades are unscanned, thus providing a relevant challenge. Our algorithm successfully classified and generated the geometry for all interior walls, with exterior walls modeled using a predefined thickness. Figure 16 illustrates the resulting IFC model, with the upper ceiling slab hidden to enhance clarity. This model accurately reflects the volumes, positions of openings, and room dimensions. Figure 16 also shows the absolute distance from the points to the wall surfaces, typically within millimeters; a similar accuracy with the deviations up to 2.5 mm in case of walls was achieved also by (Shi et al., 2019). In contrast, the VOX2BIM+ algorithm (Martens and Blankenbach, 2023) exhibited significantly more significant deviations.
@@
The dataset also facilitated testing our histogram-based technique complemented by heuristic rules for detecting openings. All internal doors were accurately identified, although the detection was occasionally compromised by furniture or large gaps in point density on the walls. Such errors require manual corrections during model editing in any BIM software.

### v2

@@ 2.8 Identification of zones
The detection of zones is typically approached through three primary methods [43]: (i) slice-based modeling, (ii) mesh simplification modeling, and (iii) floorplan-based modeling. Once the exact positions and geometries of walls are identified, Cloud2BIM circumvents these computationally intensive methods. Instead, the identify_zones function (Figure 13) analyzes wall geometry by first processing the axes of walls to ensure connectivity and proper alignment. Wall segments are split at intersections of these axes and extended to connect with adjacent axes within a distance specified by the snapping_distance parameter. Subsequently, segments parallel to the wall axes, representing the surfaces of the walls, are generated at a distance equal to half the width of the individual walls. These parallel segments are then extended to intersect fully, forming the boundaries of zones. The function ultimately creates closed polygons defined by the wall axes; each polygon serves as a candidate zone. The complete process of zone identification is illustrated in Figure 14.
Figure 13: Flowchart diagram illustrating the process of identifying zones from a floor plan. The implementation of this workflow can be found in the space_generator.py file, function identify_zones, lines 887–928 [32].
@@ 2.9 Identification of openings
Several methods for detecting openings have been developed over the years. Techniques for identifying door openings have typically been categorized into histogram-based approaches [11, 28] and those that utilize scanner trajectory data [14]. Each method, however, comes with specific limitations. Density-based analysis can be sensitive to obstructions like furniture, affecting the histogram’s accuracy, while trajectory-based methods depend heavily on particular scanning technology. Due to these constraints, our approach utilizes a density analysis from histograms, as illustrated in the flowchart in Figure 15, where the function identify_openings is outlined.
Figure 15: Flowchart diagram illustrating the process of identifying openings from the distribution of points assigned to a wall. The implementation of this workflow can be found in the aux_functions.py file, function identify_openings, lines 1181–1331 [32].
@@
The analysis begins by positioning the wall at the coordinate origin and rotating it into the $x$-$z$ plane. The floor plan position of the opening, primarily along the $x$-axis, is first determined. The histogram is then analyzed based on a threshold defined as a percentage of the tenth maximum value recorded to identify the start and end of gaps or low-density areas. The tenth maximum was selected based on observations of the histograms and found appropriate to avoid extreme maxima that seldom occur in the histogram; in some walls there were about 1–3 extreme maxima so that the tenth maximum provides a safe reserve, but the number of selected maximum can be changed. When matching openings are detected, this procedure is repeated for the $z$-axis histogram. Figure 16 presents a side view of the wall, displaying histograms for the $x$ and $z$ axes with marked door openings in red. The effectiveness of this analysis is improved by the prior dilution of the point cloud to the minimum distance, as discussed in Section 2.1. A more uniform point distribution minimizes the risk of misjudging the thresholds for empty space in the histograms, thereby improving the accuracy of opening detection.
@@ 2.10 Creation of IFC output file
The industry foundation classes (IFC) file format is primarily designed to facilitate data transfer among various participants in the modeling process, allowing the sharing of the building information model and its associated data. Utilizing IFC as the final format to transfer data from point clouds to BIM is a well-established approach, as evidenced by similar implementations by Martens and Blankenbach [10], which included the reconstruction of slabs, floors, and walls. There are numerous scan-to-BIM methodologies discussed in the literature, such as those by Adán et al. [44] and Truong-Hong and Lindenbergh [45]; however, these approaches typically do not support direct export to IFC or any standardized BIM format. The capability to export directly to IFC, as implemented in our Cloud2BIM tool, facilitates seamless integration with a wide range of BIM software platforms. This compatibility significantly enhances the utility of the generated models for architectural planning, structural analysis, and lifecycle management, ensuring that stakeholders can leverage the full spectrum of BIM functionalities for efficient project execution and maintenance.
Our output of Cloud2BIM is generated in the IFC 4 ADD2 TC1 format using the IfcOpenShell library for Python [46]. This process includes specifying all necessary attributes for each IFC class. We chose the DesignTransferView_V1.0 Model View Definition to ensure that our models are compatible with and can be further edited in BIM software platforms such as ArchiCAD [^3] https://graphisoft.com/solutions/archicad, Revit [^4] https://www.autodesk.com/products/revit, and Allplan [^5] https://www.allplan.com.
@@ 2.11 Inputs and calibration parameters
The calibration parameters in this study were selected based on empirical testing across multiple datasets to balance accuracy and computational efficiency. As Cloud2BIM evolves, a graphical user interface (GUI) will enable real-time adjustment of all parameters with a live preview of segmentation results, ensuring flexibility for diverse datasets.
For slab detection, dilation_meters and erosion_meters control morphological operations, compensating for occlusions such as furniture covering floors. These values must be fine-tuned based on dataset conditions. The smoothing_factor, regulating Douglas-Peucker simplification, remains minimal compared to overall element dimensions and typical point cloud measurement uncertainties. The z_step parameter, chosen heuristically, prevents peak loss in histograms while maintaining distinct slab surface separation, and max_n_points_array controls point density thresholds in slab detection.
For wall segmentation, threshold ensures all nonzero values contribute to mask generation, while square(5) determines kernel size in morphological processing. The angle_tolerance parameter, heuristically set, governs collinear segment merging. In opening detection, max10 eliminates extreme peak values in histograms, with 10 chosen heuristically to filter out statistical outliers, though higher values could be considered.
@@ 3.1.1 Non-commercial tools
Multi-storya
Slabsb
Non-Manhattan wallsc
Wall surfacesd
Volumetric wallse
Volumetric spacesf
Openingsg
IFC outputh
Open-sourcei
Wang et al. [9]
Macher et al. [27]
Jung et al. [20]
Shi et al. [37]
Ochmann et al. [18]
Bassier and Vergauwen [19]
Romero-Jarén and Arranz [8]
Gen and Gentes [47]
Martens and Blankenbach [10]
Mahmoud et al. [11]
Mehranfar et al. [48]
@@
a Automatic processing of multi-story buildings, where floors are detected and structured hierarchically (e.g., as IfcBuildingStorey) without requiring manual intervention.
b Detection of slab position, thickness, and horizontal dimensions.
c Support for walls that are not strictly aligned to an orthogonal (Manhattan) grid, allowing for arbitrary angles.
d Identification and semantic segmentation of wall surfaces from point clouds.
e Detection of walls as solid objects, including their topology (connections between walls) and representation as parametric elements with thickness, start and end points, and height.
f Creation of volumetric room representations, which are derived separately from wall geometry and are not necessarily included in solutions that detect volumetric walls.
g Geometric representation and placement of openings within wall elements, including width, height, position, and type (door, window).
h Automatic conversion of extracted building geometry into the IFC format without reliance on proprietary software.
i The tool’s code is publicly available, allowing free use and modification.
@@
In reviewing the contributions of others, our discussion emphasizes the distinction in methodologies, scope, and outcomes. We highlight how different authors have approached the segmentation of point clouds and the geometric reconstruction of architectural elements, pointing out the strengths and limitations of their approaches compared to ours. For instance, Bassier and Vergauwen [19] focuses solely on walls, creating geometry in Revit through a Rhino plug-in and then exporting to IFC, including zones. Our method supports direct IFC export from Python, reducing dependence on commercial software. In addition, their approach, like some others, lacks the ability to detect and model doors and windows, often resulting in a lower level of detail. This study though, unlike the other, allows
@@
The work of Mahmoud et al. [11] incorporates advanced deep learning techniques for semantic segmentation, followed by RANSAC for line extraction to identify architectural features. While it provides detailed segmentation of walls, floors, and ceilings, the geometry export is limited to an Excel format, which is then used by a Dynamo add-in for Revit, constraining the output to specific commercial software. Similarly, the work by Mehranfar et al. [30] relies on the Revit API for 3D geometry and employs the PointTransformer network with DBScan clustering, also utilized in an improved version by Czerniawski et al. [49]. However, this approach is sensitive to auxiliary parameter settings and does not enhance functionality beyond our method. Furthermore, it only supports single-floor point clouds, limiting its utility for multi-storey buildings.
@@
Martens and Blankenbach [10] presented VOX2BIM+, a robust approach by analyzing voxels to improve noise resistance and accuracy in the volumetric reconstruction of spaces and structural elements. Unlike our approach, which utilizes direct point cloud processing, their method incorporates several advanced techniques, including region growing [50] and morphological skeletons for wall detection. However, their code remains proprietary, and their method does not support window or door geometry extraction.
@@
Similarly, Wang et al. [9] and Ochmann et al. [18] have made significant contributions to the field, focusing on energy simulation models to detailed wall segmentation. These studies typically rely on simplified models or specific algorithms like RANSAC, which may only capture some architectural details such as doors and windows, a capability our software aims to address comprehensively. While sufficient for their goals, the detected zones that are clustered according to their purpose and not published to IFC cannot be further edited and exported to other software.
3.1.2 Commercial tools
To provide additional context on commercial solutions in the point-cloud-to-BIM domain, Table 8 summarizes key functionalities of various tools. The table highlights the presence of semantic segmentation capabilities, the ability to detect walls and openings, multi-story support, IFC output, and integration with Autodesk Revit.
Table 8: Overview of commercial tools for point cloud to BIM conversion, summarizing their key functionalities.
Tool Semantic segmentation Walls and openings detection Multi-story support IFC output
Pointfuse/Autodesk Recap
Aurivus
Faramoon a
usBIM.Scan2IFC b
Undet c
ClearEdge3D EdgeWise d
Geo-Plus VisionBIM
BricsCAD BIM v24.0 e
Leica CloudWorx f
a Exports to IFC at LOD100, requiring manual refinement.
b Multi-story support is available, but users must manually define height levels.
c Wall and opening detection requires manual selection (Fit function).
d Walls are detected, but openings are not. Wall detection accuracy is limited.
e Wall detection is available, but openings are not detected.
f Wall and opening placement is assisted but requires manual selection.
Several commercial solutions focus on semantic segmentation of point clouds, identifying elements such as walls, floors, and openings (e.g., Pointfuse, Aurivus, Faramoon, and Geo-Plus VisionBIM). However, their approach to BIM conversion varies significantly. Pointfuse / Autodesk Recap provides semantic segmentation and basic geometric modeling but lacks robust multi-story automation. Aurivus specializes in semantic classification of point clouds without generating BIM models directly, instead offering a Revit plugin for further modeling. Faramoon generates an initial IFC model at LOD100, which is then manually refined. Geo-Plus VisionBIM automates part of the modeling process, including AI-assisted window detection, but still requires substantial user input to define structural parameters.
IFC export is available in Pointfuse, Faramoon, usBIM.Scan2IFC, and Geo-Plus VisionBIM. However, many solutions do not generate fully structured BIM elements and instead rely on user intervention within proprietary software. usBIM.Scan2IFC offers multi-story support, but the user must manually define height levels, limiting automation.
Faramoon operates as a black-box service, where users submit point clouds, and the company returns a modeled BIM representation. This contrasts with tools like Undet and Leica CloudWorx, which function as Revit plugins that assist users in manually aligning and placing elements rather than offering full automation. These tools employ a fit-based approach, where the user clicks on areas where elements should be placed, and the software adjusts the positioning accordingly.
ClearEdge3D EdgeWise and BricsCAD BIM v24.0 incorporate additional automation features. However, ClearEdge3D EdgeWise struggles with inaccurate wall connections and lacks opening detection, making its outputs less consistent. Multi-story support remains limited across most solutions, often requiring floor-by-floor processing. While BricsCAD BIM v24.0 allows users to classify point clouds and generate zones, it does not detect openings or support fully automated multi-story modeling. In contrast, Geo-Plus VisionBIM does support multi-story workflows, though with significant manual input required to refine the detected elements.
@@ 3.2 Precision analysis
This section evaluates the precision of the segmentation, classification, and reconstruction algorithms implemented in Cloud2BIM using the datasets introduced in Section 2.1. Although established metrics exist for quantifying the precision of 3D model reconstructions [51], these were not employed in our evaluation due to their dependency on manually created ground-truth models, which were not available. Instead, we present visual assessments complemented by heat maps illustrating absolute distances between the original point clouds and the resulting IFC model surfaces. These visualizations effectively demonstrate the achieved precision of the proposed methods across all tested datasets.
@@ 3.2.1 ETH Zürich synth3
The ETH Zürich synth3 dataset was selected to showcase the performance of our volumetric method in wall segmentation within a 2.5-dimensional Atlanta-world system. This dataset is frequently used to benchmark the success of interior wall segmentation [18, 19, 37]. Although it lacks an exterior scan, this limitation mirrors real-life scenarios where facades are unscanned, thus providing a relevant challenge. Our algorithm successfully classified and generated the geometry for all interior walls, with exterior walls modeled using a predefined thickness. Figure 18 illustrates the resulting IFC model, with the upper ceiling slab hidden to enhance clarity. This model accurately reflects the volumes, positions of openings, and room dimensions. Figure 18 also shows the absolute distance from the points to the wall surfaces, typically within millimeters; a similar accuracy with the deviations up to 2.5 mm in case of walls was achieved also by [37]. In contrast, the VOX2BIM+ algorithm [10] exhibited significantly more significant deviations.
@@
The dataset also facilitated testing our histogram-based technique complemented by heuristic rules for detecting openings. All internal doors were accurately identified (13/13), although the detection was occasionally compromised by furniture or large gaps in point density on the walls. Such errors require manual corrections during model editing in any BIM software (see Section 3.4, Figure 23).
@@ 3.2.3 Prague hotel
3.2.4 ISPRS benchmark datasets
Despite achieving satisfactory reconstruction results, we did not use the ISPRS benchmark datasets [52] to evaluate the accuracy of Cloud2BIM. These datasets, designed to support research in automated 3D indoor modeling, include predefined evaluation metrics based on completeness, correctness, and accuracy. However, these metrics require a manually created ground-truth model, which introduces potential inaccuracies. Furthermore, we identified several obstacles in these datasets, including occlusions, clutter, and non-standard architectural features, making them less suitable than those selected in Section 2.1 for assessing Cloud2BIM’s performance. Since we already had dedicated benchmark datasets, no additional ones were included to maintain the study’s focus. However, Cloud2BIM is freely available [32], and readers are encouraged to conduct their own evaluations using the ISPRS datasets.
To illustrate the characteristics of the ISPRS benchmark datasets, we provide an overview of individual point clouds along with figures demonstrating their reconstruction using Cloud2BIM:
  • 1.
    Fire Brigade (Delft, The Netherlands): This dataset was acquired using a terrestrial laser scanner and is characterized by a high level of clutter, as noted by the dataset authors. Structural elements are partially obstructed by furniture and other objects, and occlusions in static laser scanning result in data gaps. Despite these challenges, Cloud2BIM successfully reconstructed external walls and detected most openings (Figure 21, top image). The indoor scene contains 9 rooms, 10 doors, and 53 windows, with additional complexity introduced by curtain walls containing windows.
  • 2.
    Grainger Museum (Melbourne, Australia): This dataset represents a Non-Manhattan World building with 18 rooms, 13 doors, and 41 windows. Captured using the Zeb Revo RT system from both indoors and outdoors, it features significant clutter due to the presence of various artifacts. Curved walls posed a major challenge, as Cloud2BIM is designed for rectilinear structures. When detected, these curved walls were approximated using straight segments, highlighting a limitation of the current workflow (Figure 21, bottom image).
  • 3.
    TUB1 (Technical University of Braunschweig, Germany): Captured with the Viametris iMS3D system, this dataset includes both a point cloud and the sensor trajectory. It features 10 rooms enclosed by walls of varying thicknesses, 23 doors (open and closed), and 7 windows. As the building is unfurnished, the clutter level is low, primarily consisting of points corresponding to people present during the survey. Cloud2BIM handled this dataset well, although variable floor levels posed challenges, requiring careful tuning of input parameters (Figure 22, top image).
  • 4.
    TUB2 (Technical University of Braunschweig, Germany): Acquired in the same building as TUB1 but using a Zeb-Revo scanner, this dataset covers two floors connected by a staircase. The first level contains 14 rooms, 8 windows, and 23 doors, while the second level consists of 10 rooms, 13 windows, and 28 doors. Wall thicknesses and ceiling heights vary. Like TUB1, the dataset has low clutter, making it a relatively straightforward test case for Cloud2BIM (Figure 22, bottom image).
Figure 21: Reconstruction results for the ISPRS Fire Brigade (top) and Grainger Museum (bottom) benchmark datasets; the point clouds are color-coded by height (blue to red), while the reconstructed IFC elements, including slabs, walls, and detected openings, are displayed in gray.
Figure 22: Reconstruction results for the ISPRS TUB1 (top) and TUB2 (bottom) benchmark datasets; the point clouds are color-coded by height (blue to red), while the reconstructed IFC elements, including slabs, walls, and detected openings, are displayed in gray.